OpenStudy (anonymous):

Need help with Triangles! Please help! If you could answer all that would be great I am really having trouble with this section. Find the remaining angles and sides in each triangle. 1. In Triangle DEF , d = 24 m, e = 21 m, and f = 25 m. Simplify each trigonometric expression. 2. sin 40º cos 50º + cos 40º sin 50º 3. cos 105º cos 15º – sin 105º sin 15º 4. cos 55º cos 25º + sin 55º sin 25º 5. sin 70º cos 10º – cos 70º sin 10º

OpenStudy (anonymous):

\[d^2=e^2+f^2-2ef*\cos(\alpha)\] Solve for alpha and get the answers for number one, by switching around d, e, and f

OpenStudy (anonymous):

2, 3, 4, 5, there are identities like sin(2x), which you can simplify those expressions into.

OpenStudy (anonymous):

2. \[sin(90)=0\]

OpenStudy (anonymous):

3 \[cos(105+15)=cos(120)=1-2cos^2(60)=1-2/4=1/2\]

OpenStudy (anonymous):

I have done 2, and 3 for you, you can do 4 and 5 using the same technique